Solve the following system of linear equations by graphing. (If the system has no solution, enter NONE for each answer. If the system is dependent, enter INFINITELY MANY for each answer.)
3x+y=0
x-3y=0
x = ?
y = ?
pls show me the steps

Answer by stanbon(75887)

3x+y=0
x-3y=0
---------
Multiply thru 1st by 3 to get:
9x+3y = 0
x -3y = 0
---------
Add and solve for "x":
10x = 0
x = 0
solve for "y":
3x+y = 0
0+y = 0
y = 0
====================
x = 0
y = 0
----
Cheers,
Stan H.